mathslover (mathslover):

Application of Derivatives Find the angle of intersection between 2 curves : \(y^2 = 4x\) and \(x^2 = 4y\) .

mathslover (mathslover):

@ganeshie8

ganeshie8 (ganeshie8):

the two curves intersect at (0, 0) and (4, 4)

mathslover (mathslover):

Yeah.

ganeshie8 (ganeshie8):

angle of intersection at (0,0) is trivial - horizontal/vertical tangents - that means angle between them is 90 degrees

mathslover (mathslover):

Oh...

mathslover (mathslover):

Fine now. I can continue from here!

ganeshie8 (ganeshie8):

find the tangents at (4, 4) and find the angle between them.... or do u have any shortcut formula for this also ? :)

mathslover (mathslover):

Well, dy/dx = 4/2y (from y^2 = 4x) At (4,4) (dy/dx) = 4/8 = 1/2 = m1 and At (0,0) (dy/dx) = 4/0 = infinity Also, dy/dx = x/2 (from x^2 = 4y) At (4,4) (dy/dx) = 4/2 = 2 = m2 and at (0,0) dy/dx = 0

mathslover (mathslover):

Angle between 2 curves : \(\tan \theta = \cfrac{m_1 - m_2}{1 + m_1 m_2 }\) Theta = tan^{-1} (3/4)

mathslover (mathslover):

that is not a shortcut but yes, that is what I know yet :)

ganeshie8 (ganeshie8):

yes i would do it the exact same way :)

mathslover (mathslover):

Great then. Thanks again :)

mathslover (mathslover):

Good Night Bhaiya! :-)

ganeshie8 (ganeshie8):

this also works : http://www.wolframalpha.com/input/?i=arctan%282%29+-+arctan%281%2F2%29+

ganeshie8 (ganeshie8):

good night, have good sleep :)

mathslover (mathslover):

After getting such kind of help, I will surely have a sound sleep. :) Gn